OpenStudy (anonymous):

The function f(h) shown below represents a plumber's total charges, inclusive of call out charges, for h hours of work: f(h) = 70h + 50 Calculate and interpret the meaning of f(2).

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

f(2) = 70(2) + 50 = 140 + 50 =190